OpenStudy (anonymous):

X^2-16x+64/X^2-64 need help simplifying this ratio

OpenStudy (akashdeepdeb):

BIG HINT: In the numerator use [(a-b)^2 = a^2 - 2ab + b^2] In the denominator use [a^2-b^2 = (a+b)(a-b)] Okay? :D

OpenStudy (anonymous):

okay thanks

OpenStudy (akashdeepdeb):

:)